The speed of DNA replication at a replication fork is about 100 nucleotides per second per standin human cells. What is the minimum number of origins of replication that a human cell must have in order to replicate its DNA once every 24 hours? Recall that a human cell contains two copies of the human genome, one inherited from the mother and one from the father, each consisting of 3 x 109base pairs.
Answer #1

at an origin of replication, 2 replication forks are formed, it moves in opposite directions, so rate of DNA synthesis from an origin of replication=100\times2=200 nucleotides/s

time taken to replicate if there was only one origin of replication= length of the DNA/rate of synthesis at one origin of replication

= 6 \times 10^9/200

=3 \times 10^7 S

1 hour=60 minute

1 minute=60 second

so 1 hour=60\times60=3600S

converting 3 \times 10^7 S to hour

3 \times 10^7 S/3600=8333.34 hours

but the time of DNA replication=24 hours

so there has to be multiple origin of replication

number of origin of replication required to complete the replication in 24 hours= time taken to replicate the DNA/t24

=8333.34 hours/24

= 347.2

so the number of origin of replications=347.2